# references.bib

### v1

@@ 3.3 Validation Effectiveness
When delegating information production to human experts, there is a challenge of ”hidden action,” where the actions of experts are not fully observable or assessable, potentially leading to misalignment of interests [Gurbaxani1990]. Similarly, when delegated to AI, the issue shifts also to the transparency and reliability of outputs. While human experts may act in self-interest, leading to agency costs through false information or requiring additional incentives, AI systems present the challenge of output opacity. End users cannot evaluate the correctness of SQL queries and resulting KPIs until after implementation, risking decisions based on false information with economic consequences. Ensuring the reliability of AI-generated information is crucial, particularly in high-stakes decision-making scenarios where false information can result in significant negative business outcomes. The requirement for reliability is further intensified by the inherent tendency of LLMs to produce hallucinations [Lin2024, Perkovic2024], resulting in fabricated or erroneous SQL queries that may appear credible. Consequently, robust techniques are essential to ensure the validation of these generated SQL queries.
The techniques for validating automatically generated SQL queries fall within the broader field of explainable AI [DARPA2016, Rudin2019, Minh2022, Langer2021, Gunning2019]. Since the advent of expert systems, researchers have emphasized the pivotal role of explanations in enhancing the interpretability of AI-generated results [Wallis1982, Nakatsu2006]. In recent years, the growing demand for transparency in machine learning has prompted significant advances in the field of explainable AI, leading to an expansion in both the scope and sophistication of explanation techniques. The primary objective of explainable AI is to provide explanations that elucidate the underlying processes behind AI outputs, thereby facilitating tasks such as debugging, regulatory compliance, the establishment of user trust, and the effective utilization of AI.
In the specific context of Text-to-SQL, validation necessitates the provision of explanations that help to comprehend and verify AI-generated SQL queries. These explanations must detail the structure of the queries, including the involved tables, fields, and operations (e.g., filtering, grouping, or joining). Such detailed information enables users to critically evaluate the logical coherence of the query and its alignment with their intended objectives, thereby supporting an independent validation process. Effective validation ensures that users can reliably differentiate between correct and incorrect SQL queries. However, ineffective validation—where erroneous queries are mistakenly validated or accurate ones are erroneously dismissed—compromises the reliability of the process and may result in suboptimal decision-making.
@@
Figure 5: Decomposition technique for explaining SQL queries [Arpit2021].
@@
Current literature identifies three different techniques for validating and explaining SQL queries (see [Zhang2023__] for an overview). Decomposition techniques (see Figure 5) break SQL queries into components such as tables, fields, and operations, presenting intermediate results to trace each element’s contribution to the final output, aiding error detection [Arpit2021]. Visualization techniques use graphical representations to map relationships among query components, enhancing clarity and navigation in complex query structures [Leventidis2020, Miedema2021, Murakawa2011]. Dialogue techniques employ LLMs to provide natural language explanations, enabling users to interactively explore and refine queries by addressing specific components or the underlying data model [Tian2024, Elgohary2020, Kokkalis2012, Gur2018, Xiaxia2021, yao2019model].
@@
Empirical studies on validation techniques for Text-to-SQL reveal varying levels of effectiveness in supporting SQL comprehension and error correction. Ning et al., through a user study, observed that decomposition, visualization, and dialogue techniques achieved an effectiveness rate of approximately 56% [Zhang2023__]. In contrast, Tian et al. demonstrated that their technique, which combines decomposition and dialogue techniques, significantly outperformed existing techniques, achieving a validation and error correction effectiveness of approximately 85% [Tian2024]. These findings underscore substantial progress in user-based validation and error correction, showcasing their transformative potential to enhance the reliability of AI-generated SQL queries. However, the results also highlight persistent challenges, particularly in real-world scenarios involving large-scale, highly complex SQL queries and extensive data models, where existing techniques may prove inadequate.
A critical factor influencing the success of validation is the interplay between the competence of the end user and the quality of the explanations provided. While Tian et al. demonstrated that their technique consistently achieved high effectiveness regardless of users’ technical expertise [Tian2024], the intricacies of SQL queries in complex business environments such as data warehouses with hundreds of interconnected tables and fields necessitate a more nuanced evaluation.
The end user’s background and domain experience significantly affect their ability to assess the correctness of SQL queries, and this capability is further shaped by the clarity, detail, and comprehensiveness of the explanations generated. Additionally, the complexity of the underlying data model [Chan2005] and the resulting SQL query plays a pivotal role in the validation outcome. Complex data models with numerous interdependencies and abstract schema structures can obscure the logic of even well-constructed SQL queries, making validation tasks challenging for end users.
@@ 4.1 Basic Assumptions
This study introduces models based on rational choice theory, with an emphasis on expected utility theory [Neumann2007, Tunney2024], to evaluate the effective use of CBA. The analysis focuses on the example of Text-to-SQL. Two primary influencing factors are considered:
@@
In an ideal scenario, both accuracy and validation effectiveness would be perfect, ensuring the generation of reliable and actionable information. However, these factors are inherently probabilistic and subject to imperfections. The models examine how the interaction between accuracy and validation effectiveness influences the effectiveness of Text-to-SQL and its potential advantages over delegation to a human expert. Specifically, the models evaluate how varying levels of these factors affect the relative advantage of delegating tasks to AI, with or without validation support. By exploring this interplay, the models aim to identify the conditions under which Text-to-SQL can serve as a viable alternative to human expertise, providing a structured framework for its adoption in CBA. This structured approach not only quantifies the trade-offs between accuracy and validation effectiveness but also highlights how CBA bridges the competency gaps that hinder traditional self-service analytics. By doing so, it underscores AI’s role in empowering end users to make informed decisions without relying on advanced technical skills, thereby expanding the accessibility and effectiveness of analytics.
@@
Currently, no data engineer is available to create the KPI, and such a resource will only be accessible at a later time. The delayed delivery of the KPI results in a reduced (net) profit $v$ (where $10<v<$ 1). This profit is lower than the achievable maximum of $+1$, reflecting the adverse effects of delayed action. In extreme cases, $v$ may even approach zero, rendering the KPI almost valueless. Additional factors, such as the effort required to align requirements and potential motivational challenges that could negatively impact collaboration between end user and data engineer, are not considered in this analysis. Consequently, $v$ serves as a measure of the urgency of KPI procurement, with lower values of $v$ indicating greater urgency.
The process of KPI generation can be delegated to AI equipped with a natural language interface powered by LLMs. AI offers two levels of assistance: partial support (PS) and full support (FS). Under PS, the AI translates the natural language request into SQL code, which is executed on the data warehouse to calculate the KPI. The KPI is then used to inform decision-making, resulting in a business impact. FS includes an additional validation step to ensure the correctness of the SQL query. During this step, the end user is provided with explanations of the generated SQL query, enabling independent validation of its correctness. If the SQL query is deemed correct, the process concludes with confirmation, allowing to act on the KPI. If the query is identified as erroneous, it is rejected, and no action is taken, avoiding both gains and losses. In this process, AI acts as a collaborative partner, assisting in overcoming technical challenges while ensuring that the business decision is supported by a correct and validated KPI.
@@
Since traditional self-service analytics is not feasible due to a lack of expertise, the end user evaluates whether the KPI should be generated via delegation to AI (either PS or FS), or whether it is more advantageous to wait for the data engineer to receive the KPI at a later time. The end user selects the option with the highest benefit. When CBA (either PS or FS) offers a higher expected value than the profit through the data engineer, it is preferred. Conversely, if the expected value from CBA is lower, human delegation is preferred. In such cases, CBA is deemed ineffective, as it fails to provide timely and decision-relevant information. By addressing the lack of technical skills required for traditional self-service analytics, CBA closes the competency gap that often hinders non-technical users from independently accessing and analyzing data.
@@ 4.3 Full Support (FS)
The second stage involves assessing the viability of FS, which includes an additional validation process to enhance the reliability of AI-generated SQL queries. This evaluation determines whether validation meaningfully improves the quality of the generated queries, ensuring their correctness and relevance for decision-making. By incorporating validation, FS demonstrates AI’s role as an active partner in ensuring reliability and correctness. The validation process not only identifies and mitigates errors but also enhances the end user’s trust in the system by providing transparent explanations. This collaboration between AI and the end user enables decision-making that aligns more closely with organizational goals, even in scenarios with imperfect accuracy.
Various techniques such as decomposition, visualization, and dialog-based are used to enhance the interpretability and traceability of automatically generated SQL queries. The effectiveness of these validation techniques depends on several factors, including the clarity, comprehensiveness, and contextual relevance of the explanations provided, as well as the user’s technical expertise in evaluating them. High-quality explanations can significantly aid users in identifying errors and verifying the correctness of SQL queries. However, the end user’s ability to critically evaluate these explanations plays an equally crucial role in determining the success of the validation process.
Validation is inherently uncertain due to variability in user comprehension and the probabilistic nature of AI-generated explanations. These explanations are not guaranteed to be correct, as LLMs generate outputs based on probabilities rather than deterministic logic. Poor validation can amplify risks by failing to detect erroneous SQL queries or by incorrectly rejecting valid queries. This dual risk undermines decision-making and may reduce trust in the system. While user-based validation is intended to mitigate inaccuracies, its effectiveness heavily depends on the user’s ability to interpret complex SQL logic and explanations. This reliance introduces variability, making validation inconsistent across users with differing expertise levels.
To quantify this uncertainty, validation effectiveness is modeled as a random variable, denoted as $\beta$ (where $10<\beta<$ 1). This probability reflects the likelihood that a user can correctly assess the SQL query based on the explanations provided. It is defined as the ratio of successful validations to the total number of validation attempts. By adopting this probabilistic framework, the model accounts for the inherent uncertainties associated with validation processes.
The expected value $E_{FS}(\alpha,\beta)$ for FS is calculated as:
@@ 4.4 Comparison of Full Support with Partial Support and Delegation to the Data Engineer
This condition (2) requires the validation to be sufficiently effective and robust (e.g., against the shortcomings of the end user conducting the validation based on the provided explanation) to compensate for inaccuracies in AI-generated information (($1-\alpha$)) and to deliver an overall expected value exceeding the guaranteed, but delayed profit $v$. If the AI delegation condition for FS (2) is satisfied, it is prioritized over delegation to the data engineer. However, this preference also depends on the accuracy of AI-generated information being sufficiently high compared to the profit $v$. Specifically, the ”FS feasibility condition” must hold:
@@
In condition (4), $\beta^{**}$ represents the threshold where validation effectiveness exceeds accuracy, making FS more advantageous than PS. It is important to note that all conditions described so far strongly depend on the values of the potential gain ($+1$) and loss ($-1$) and the assumption that their magnitudes are equal. If the value of the loss outweighs that of the gain, the threshold for $\beta$ in the validation dominance condition shifts, allowing a less stringent validation standard for FS to outperform PS.
@@
In region $A$, the FS feasibility condition (3) is not met, as the accuracy of information generation is less than or equal to the profit achievable through the data engineer ($\alpha\leq v$). In this situation, neither PS nor FS offers sufficient expected value. PS fails due to low accuracy, and FS is unable to compensate for this limitation, even when validation effectiveness would be perfect. Delegation to the data engineer remains the only viable option for retrieving the KPI. In this scenario, CBA is not utilized because it fails to match the effectiveness of the data engineer, rendering CBA unsuccessful. Furthermore, if the data engineer provides the KPI with significant delay, rendering it irrelevant for the end user due to its diminished value for timely decision-making ($v$ approaching zero), this not only signifies the failure of CBA but also highlights a broader limitation of business analytics in fulfilling its fundamental objective: delivering timely and decision-relevant information.
@@
In Region $B$, the FS feasibility condition (3) is met ($\alpha>v$), but the AI delegation condition for PS (1) is not satisfied ($\alpha\leq\alpha^{*}_{PS}$). In this case, PS is less effective than delegating to the data engineer. However, FS becomes a viable option if validation effectiveness meets the AI delegation condition for FS (2). Under this condition, validation introduces a ”boosting effect” by identifying and mitigating errors in SQL queries, ensuring that flawed KPIs do not influence the decision-making process. This boosting effect enables FS to enhance the overall reliability of information, compensating for the limitations of lower accuracy. For example, even if accuracy falls below the threshold required for PS (1) (e.g., $\alpha_{1}$), sufficiently high validation effectiveness in FS (e.g., $\beta_{1}$) can still yield an expected value greater than $v$. While the generation of the SQL query may fail, resulting in incorrect KPI, sufficiently effective validation can still identify these errors, ensuring that incorrect KPIs do not adversely affect the decision process.
In region $C$, where the AI delegation condition for PS (1) is met ($\alpha>\alpha^{*}_{PS}$), PS becomes a viable alternative to delegation to the data engineer, as it provides sufficient utility on its own. If the AI delegation condition for FS (2) and the validation dominance condition (4) are not fulfilled (e.g., at $\alpha_{2}$ and $\beta_{2}$), FS has a ”devastating impact” as the validation process undermines the reliability of the SQL queries generated. Not only are erroneous SQL queries less effectively identified, but insufficient validation also leads to the unwarranted rejection of correct SQL queries. Consequently, FS is not only less effective than PS but also inferior to delegation to the data engineer. Therefore, using AI-based assistance for KPI generation with AI-supported validation is not recommended. This underscores the inherent risks of relying on user-based validation without robust mechanisms to support in correctly evaluating AI-generated SQL queries. In cases where expertise is low or explanations are unclear, FS may inadvertently introduce errors or misjudgments. This highlights the critical need for improved validation tools that reduce user dependency or provide guided assistance during the evaluation process.
FS becomes favorable when, in region $C$, both conditions, (2) and (4), are fulfilled (e.g., at $\beta_{2}$ and $\alpha_{3}$). In such cases, the validation process enhances the reliability of the generated KPI and maximizes the benefits of FS. As accuracy increases, the requirements for validation effectiveness must also rise to ensure that FS remains beneficial. A particular challenge arises when accuracy is very high; in such scenarios, user-based validation must become even more effective to provide added value over PS. This places significant demands on validation mechanisms, especially in terms of users’ ability to evaluate and assess complex SQL queries based on the given explanations.

### v2

@@ 3.3 Validation Effectiveness
When delegating information production to human experts, there is a challenge of ”hidden action,” where the actions of experts are not fully observable or assessable, potentially leading to misalignment of interests [27]. Similarly, when delegated to AI, the issue shifts also to the transparency and reliability of outputs. While human experts may act in self-interest, leading to agency costs through false information or requiring additional incentives, AI systems present the challenge of output opacity. End users cannot evaluate the correctness of SQL queries and resulting KPIs until after implementation, risking decisions based on false information with economic consequences. Ensuring the reliability of AI-generated information is crucial, particularly in high-stakes decision-making scenarios where false information can result in significant negative business outcomes. The requirement for reliability is further intensified by the inherent tendency of LLMs to produce hallucinations [43, 63], resulting in fabricated or erroneous SQL queries that may appear credible. Consequently, robust techniques are essential to ensure the validation of these generated SQL queries.
The techniques for validating automatically generated SQL queries fall within the broader field of explainable AI [15, 67, 49, 37, 24]. Since the advent of expert systems, researchers have emphasized the pivotal role of explanations in enhancing the interpretability of AI-generated results [83, 55]. In recent years, the growing demand for transparency in machine learning has prompted significant advances in the field of explainable AI, leading to an expansion in both the scope and sophistication of explanation techniques. The primary objective of explainable AI is to provide explanations that elucidate the underlying processes behind AI outputs, thereby facilitating tasks such as debugging, regulatory compliance, the establishment of user trust, and the effective utilization of AI.
In the specific context of Text-to-SQL, validation necessitates the provision of explanations that help to comprehend and verify AI-generated SQL queries. These explanations must detail the structure of the queries, including the involved tables, fields, and operations (e.g., filtering, grouping, or joining). Such detailed information enables users to critically evaluate the logical coherence of the query and its alignment with their intended objectives, thereby supporting an independent validation process. Effective validation ensures that users can reliably differentiate between correct and incorrect SQL queries. However, validation is ineffective if incorrect SQL queries are misclassified as correct, or correct ones are misclassified as incorrect.
@@
Figure 5: Decomposition technique for explaining SQL queries [57].
@@
Current literature identifies three different techniques for validating and explaining SQL queries (see [60] for an overview). Decomposition techniques (see Figure 5) break SQL queries into components such as tables, fields, and operations, presenting intermediate results based on sample data to trace each element’s contribution to the final output [57]. Visualization techniques use graphical representations to map relationships among query components, enhancing clarity and navigation in complex query structures [39, 48, 53]. Dialogue techniques employ LLMs to provide natural language explanations, enabling users to interactively explore and refine queries by addressing specific components or the underlying data model [75, 20, 36, 26, 84, 90].
@@
Empirical studies on validation techniques for Text-to-SQL reveal varying levels of effectiveness in supporting SQL comprehension and error correction. Ning et al., through a user study, observed that decomposition, visualization, and dialogue techniques achieved an effectiveness rate of approximately 56% [60]. In contrast, Tian et al. demonstrated that their technique, which combines aspects of decomposition and dialogue techniques, significantly outperformed existing techniques, achieving a validation and error correction effectiveness of approximately 85% [75]. These findings underscore substantial progress in user-based validation and error correction, showcasing their transformative potential to enhance the reliability of AI-generated SQL queries.
Tian et al. have shown that their technique consistently delivers effective results regardless of the end users’ technical competency [75]. However, the boundaries of user-based validation in complex business contexts, such as in data warehouses with hundreds of interrelated tables and fields, require further research. The expertise and experience of the end users play a central role in their understanding of the software code and, consequently, in their ability to evaluate the correctness of the generated information. The explanatory information in this context differs significantly from explainable AI used to interpret machine learning models. Explanations in machine learning are typically tied directly to business language, such as: "The model predicts that the customer will have poor creditworthiness because the customer’s features exhibit these values." In contrast, the explanatory information in Text-to-SQL has a distinct technical focus, aiming to enhance the user’s understanding of how the underlying SQL query functions.
Furthermore, the complexity of SQL queries, particularly those involving numerous subqueries and joins, combined with the cognitive load they impose on users, affects their ability to comprehend the software code (see, for example, [52], which references developers rather than end users). This cognitive load is further exacerbated by the architecture of the underlying data model [8], especially in cases where data models include extensive dependencies and abstract schema structures that pose significant challenges for end users.
Additionally, end users typically perform validation only occasionally when generating new SQL queries to meet a specific information need. Consequently, they are unable to develop long-term knowledge of SQL or gain a deeper understanding of the underlying data model. This issue has also been observed in traditional self-service analytics [82], where users were unable to build experiential knowledge of the tools and data models employed for individual information generation.
@@ 4.1 Basic Assumptions
This study introduces theoretical models based on rational choice theory, with an emphasis on expected utility theory [81, 78], to evaluate the effective use of CBA. The analysis focuses on the example of Text-to-SQL. Two primary influencing factors are considered:
@@
In an ideal scenario, both accuracy and validation effectiveness would be perfect, ensuring the generation of reliable and actionable information. However, these factors are inherently probabilistic and subject to imperfections. The models examine how the interaction between accuracy and validation effectiveness influences the effectiveness of Text-to-SQL and its potential advantages over delegation to a human expert. Specifically, the models evaluate how varying levels of these factors affect the relative advantage of delegating tasks to AI, with or without validation support. By exploring this interplay, the models aim to identify the conditions under which Text-to-SQL can serve as a viable alternative to human expertise, providing a structured framework for its adoption in CBA.
@@
Currently, no data engineer is available to create the KPI, and such a resource will only be accessible at a later time. The delayed delivery of the KPI results in a reduced (net) profit $v$ (where $10<v<$ 1). This profit is lower than the achievable maximum of $+1$, reflecting the adverse effects of delayed action. In extreme cases, $v$ may even approach zero, rendering the KPI almost valueless. Additional factors, such as the effort required to align requirements and potential motivational challenges that could negatively impact collaboration between end user and data engineer, are not considered. Consequently, $v$ serves as a measure of the urgency of KPI procurement, with lower values of $v$ indicating greater urgency.
The process of KPI generation can be delegated to AI equipped with a natural language interface powered by LLMs. AI offers two levels of assistance: partial support (PS) and full support (FS). Under PS, the AI translates the natural language request into SQL code, which is executed on the data warehouse to calculate the KPI. The KPI is then used to inform decision-making, resulting in a business impact. FS includes an additional validation step to ensure the correctness of the SQL query. During this step, the end user is provided with explanations of the generated SQL query, which are designed to ensure transparency and enable independent validation of its correctness. If the SQL query is deemed correct, the process concludes with confirmation, allowing to act on the KPI. If the query is identified as erroneous, it is rejected, and no action is taken, avoiding both gain and loss.
@@
Since traditional self-service analytics is not feasible due to a lack of expertise, the end user evaluates whether the KPI should be generated via delegation to AI (either PS or FS), or whether it is more advantageous to wait for the data engineer to receive the KPI at a later time. The end user selects the option with the highest benefit. When CBA (either PS or FS) offers a higher expected value than the profit through the data engineer, it is preferred. Conversely, if the expected value from CBA is lower, human delegation is preferred. In such cases, CBA is deemed ineffective, as it fails to provide timely and decision-relevant information.
@@ 4.3 Full Support (FS)
In the second phase, the AI-generated SQL query is reviewed by the user to determine whether a correct KPI has been generated. Techniques such as decomposition, visualization, and dialog-based interaction are currently being explored to support users in comprehending the SQL query. This also aims to achieve transparency with regard to the generation of information. However, user-based validation also entails uncertainty. Both the user’s comprehension and the complexity of the SQL query and its underlying data model significantly influence the success of validation – specifically, the ability to correctly identify valid and erroneous SQL queries. Additionally, an inherent uncertainty arises in dialog-based interactions because AI-generated explanations are based on probabilities rather than deterministic rules.
To address this uncertainty, validation effectiveness is modeled as a random variable denoted as $\beta$ (where $10<\beta<$ 1). This probability reflects the likelihood that a user can correctly assess the SQL query based on the support provided. It is defined as the ratio of successful validations to the total number of validation attempts. By adopting this probabilistic framework, the model accounts for the inherent uncertainties associated with validation processes.
The end user only takes action if an SQL query is classified as correct. The probability of correctly classifying a correct SQL query is $\alpha\cdot\beta$, resulting in a gain of $+1$. Conversely, the probability of incorrectly classifying an incorrect SQL query as correct is $(1-\alpha)\cdot(1-\beta)$, resulting in a loss of $-1$. However, if a correct SQL query is mistakenly classified as incorrect, or an incorrect SQL query is correctly classified as incorrect, no action is taken, and neither a gain nor a loss occurs.
This, the expected value $E_{FS}(\alpha,\beta)$ for FS is given by:
@@ 4.4 Comparison of Full Support with Partial Support and Delegation to the Data Engineer
This condition (2) requires the validation to be sufficiently effective and robust (e.g., against the shortcomings of the end user conducting the validation based on the provided explanation) to compensate for inaccuracies in AI-generated information (($1-\alpha$)) and to deliver an overall expected value exceeding the guaranteed profit $v$. If the AI delegation condition for FS (2) is satisfied, it is prioritized over delegation to the data engineer. However, this preference also depends on the accuracy of AI-generated information being sufficiently high compared to the profit $v$. Specifically, the ”FS feasibility condition” must hold:
@@
In condition (4), $\beta^{**}$ represents the threshold where validation effectiveness exceeds accuracy, making FS more advantageous than PS. It is important to note that all conditions described so far depend on the values of the potential gain ($+1$) and loss ($-1$). If the value of the loss outweighs that of the gain, the threshold for $\beta$ in the validation dominance condition shifts, allowing a less stringent validation standard for FS to outperform PS.
@@
In region $A$, the FS feasibility condition (3) is not met, as the accuracy of information generation is less than or equal to the profit achievable through the data engineer ($\alpha\leq v$). In this region, neither PS nor FS offers sufficient expected value. PS fails due to low accuracy, and FS is unable to compensate for this limitation, even when validation effectiveness would be perfect. Delegation to the data engineer remains the only viable option for retrieving the KPI. In this scenario, CBA is not utilized because it fails to match the effectiveness of the data engineer, rendering CBA unsuccessful. Furthermore, if the data engineer provides the KPI with significant delay, rendering it irrelevant for the end user due to its diminished value for timely decision-making ($v$ approaching zero), this not only signifies the failure of CBA but also highlights a broader limitation of business analytics in fulfilling its fundamental objective: delivering timely and decision-relevant information.
@@
In Region $B$, the FS feasibility condition (3) is met ($\alpha>v$), but the AI delegation condition for PS (1) is not satisfied ($\alpha\leq\alpha^{*}_{PS}$). In this region, PS is less effective than delegating to the data engineer. However, FS becomes a viable option if validation effectiveness meets the AI delegation condition for FS (2). Under this condition, validation introduces a ”boosting effect” by identifying and mitigating errors in SQL queries, ensuring that flawed KPIs do not influence the decision-making process. This boosting effect enables FS to enhance the overall reliability of information, compensating for the limitations of lower accuracy. For example, even if accuracy falls below the threshold required for PS (1) (e.g., $\alpha_{1}$), sufficiently high validation effectiveness in FS (e.g., $\beta_{1}$) can still yield an expected value greater than $v$. While the generation of the SQL query may fail, resulting in incorrect KPI, sufficiently effective validation can still identify these errors, ensuring that incorrect KPIs do not adversely affect the decision process.
In region $C$, where the AI delegation condition for PS (1) is met ($\alpha>\alpha^{*}_{PS}$), PS becomes a viable alternative to delegation to the data engineer, as it provides sufficient benefit on its own. If the AI delegation condition for FS (2) and the validation dominance condition (4) are not fulfilled (e.g., at $\alpha_{2}$ and $\beta_{2}$), FS has a ”devastating effect” as the validation process undermines the reliability of the SQL queries generated. Not only are erroneous SQL queries less effectively identified, but insufficient validation also leads to the unwarranted rejection of correct SQL queries. Consequently, FS is not only less effective than PS but also inferior to delegation to the data engineer. Therefore, using AI-based assistance for KPI generation with AI-supported validation is not recommended. This underscores the inherent risks of relying on user-based validation without robust mechanisms to support in correctly evaluating AI-generated SQL queries. In cases where expertise is low or explanations are unclear, FS may inadvertently introduce errors or misjudgments.
FS becomes favorable when, in region $C$, both conditions, (2) and (4), are fulfilled (e.g., at $\beta_{2}$ and $\alpha_{3}$). In such cases, the validation process enhances the reliability of the generated KPI and maximizes the benefits of FS. As accuracy increases, the requirements for validation effectiveness must also rise to ensure that FS remains beneficial. A particular challenge arises when accuracy is very high; in such cases, user-based validation must become even more effective to provide added value over PS. This places significant demands on validation mechanisms, especially in terms of users’ ability to evaluate and assess complex SQL queries based on the given explanations.